Two cyclist leave a city at the same time one going east and the other going west. The west bound cyclist bike at 9 mph faster than the east bound cyclist. After 5 hour they are 225 miles apart. How fast is each cyclist riding
Answer by mananth(16949)

east bound - speed = x mph
west bound speed = x+9 mph
..
eastbound distance traveled = 5x
westbound distance traveled = 5(x+9)
...
5x+5(x+9)=225
5x+5x+45=225
10x=225-45
10x=180
/10
x= 180/10
x= 18 mph eastbound speed
x+9 = 18+9 = 27 mph westbound speed